Aiden put two-fifths of his money into his piggy bank. He had 15 dollars left in his pocket to buy a toy. How much money did Aiden have before he put money in his bank?
Answer by checkley77(12844)

Sorry I used the wrong fraction.
He put 2/5 in the bank which left 3/5 for the toy.
3/5x=15
3x/5=15
3x=15*5
3x=75
x=75/3
x=$25 ans. for the original amount.